OpenStudy (anonymous):

can somebody please tell me how to square root 63/7 and how they got to that answer?

OpenStudy (anonymous):

63/7 = 9

OpenStudy (anonymous):

\[\sqrt{\frac{ 63 }{ 7 }}=\frac{ \sqrt{63} }{\sqrt{7} }\]

OpenStudy (anonymous):

it was 3 but thanks anyway

OpenStudy (anonymous):

@bowiefan If you simplify his answer further you'll arrive at 3.

OpenStudy (anonymous):

@bowiefan yea i was getting at that but ok if u understand.